# Synthesising Activity Participations and Scheduling with Deep Generative Machine Learning

### v2

@@ 4.2.1 Discrete Embedding and Un-Embedding
Un-embedding reverses the operation using soft-max and then sampling the highest value at each time step. Note that in all cases, we find sampling the highest probabilities at each step outperforms weighted sampling.
@@ 4.2.2 Continuous Embedding and Un-Embedding
The continuous encoding is un-embedded by the decoder following a similar process in reverse. A vector of durations for the sequence is separated and passed through a sigmoid activation so that all duration values are between zero and one. The remaining vectors are decoded into activities using a soft-max layer and sampling the highest value at each step. Activities and their durations for each sequence are then concatenated back together. After discarding padding tokens, durations for each sequence are normalised such that they sum to one.
@@ 4.3 Reconstruction Losses
The discrete embedding uses the mean cross-entropy of all predictions as reconstruction loss. The final loss function for the discrete encoding-based models is therefore:
$\text{Loss}=\text{CrossEntropy}_{\text{activities}}+\beta\text{KLD}$
The continuous encoding uses a combination of mean cross-entropy loss for the activity prediction and mean squared error for the duration prediction. We introduce an additional hyperparameter $\alpha$, detailed in Table 4, to weight the duration component of loss.
$\text{Loss}=\text{CrossEntropy}_{\text{activities}}+\alpha\text{MSE}_{\text{durations}}+\beta\text{KLD}$
@@ 4.4 Encoder/Decoder Block Architectures
The following section overviews the Feed-Forward, Convolutional and Recurrent VAE architectures. The number of blocks ($N$) and their size ($S$) vary for each model and are summarised in Table 4. The full architectures are available from Caveat [^4] https://github.com/big-ucl/caveat.
@@ 4.4.3 Recurrent (RNN) VAEs
The Recurrent (RNN) model architecture is presented in Figure  LABEL:fig:rnn_template. The encoder layer uses $N$ stacked Long-Short Term Memory (LSTM) units of size $N$ (Hochreiter and Schmidhuber, 1997). The block iteratively progresses through the input embeddings, and each LSTM unit passes a learnt hidden state as input to the next. The final output hidden state is then flattened as required for the latent block.
@@ 4.5 Model Training and Hyperparameters
Models are trained on 90% of the real sample data. We use the remaining 10% for validation during training. We train models until validation loss stabilises, typically for around 100 epochs. We do not find models suffer from over-fitting, likely due to the VAE design, which adds a stochastic process to training. We use Adam for gradient descent. Model training hyperparameters are reported in Table 4. The hyperparameters are chosen to optimise validation loss using the Tree-structured Parzen Estimator by Optuna (Akiba et al., 2019).
Table 4: Models Overview
@@ 5 Evaluation Methodology
In this section, we consider the evaluation of the synthetic samples of activity schedules $\hat{X}$. Our aim is for the synthetic distribution $\hat{P}_{X}$ to be equal to the real distribution $P_{X}$. This represents the ability of a model to correctly estimate the probability of any possible activity schedule - commonly called density estimation. We estimate both $P_{X}$ and $\hat{P}_{X}$ by extracting marginal distributions and comparing these between the real and synthetic samples using a distance metric. We choose marginal distributions with qualitative meaning such that we can assess the implications of the quality of density estimation on specific tasks. We consider many marginal distributions such that when distances are aggregated, we can make a reasonable quantitative evaluation of overall density estimation. Density estimation is detailed in Section 5.1.
Perfect density estimation would satisfy our requirement for both individual schedule quality and aggregate representation across multiple schedules. However, because we can only estimate densities and because perfect estimation is not feasible, we supplement evaluation by additionally considering metrics for individual sample quality (Section 5.2) and creativity (Section 5.3).
@@ 5.1.1 Marginal Distributions
Both participation and transition distributions are measured as rates, which are counts of occurrences for each schedule. We then consider the distribution of such rates across the whole sample. The left-hand side of Figure 10 shows an illustrative example of rate distributions for a real and synthetic sample of schedules.
@@
Figure 10 illustrates rate (left-hand-side) and time (right-hand-side) marginal distributions. Rates are ordinal, and timings are continuous. However, the time distributions are binned at five-minute intervals to reduce computational complexity. All distributions are described using mean density, which can be interpreted as the expected rate or time.
@@ 5.1.2 Distances
Distances between the real and synthetic distributions are measured using Earth Movers Distance (EMD). EMD is symmetric with meaningful values, such that we can reasonably interpret them. For example, a distance of 0.1 for work activity participation rate can be interpreted as an expectation of 0.1 fewer or more of work activities occurring in each schedule. Similarly, a distance of 0.5 for an activity start can be interpreted as that activity starting, on average, half a day late or early. For bivariate distributions, we use L1 distance (also known as Manhattan/city-block distance).
Figure 10: Illustrative Marginal Distributions
@@
Domain Marginal Distribution Units Descriptive Metric
Description Segmentation
@@ 5.3 Creativity Evaluation
We are additionally interested in the model’s ability to generate both (i) a good diversity of schedules – i.e. different from one another, as well as (ii) novelschedules – i.e. unique from schedules in the real sample. Diversity can be considered as a type of density evaluation metric, but is also used to specifically check if the model is suffering from mode collapse, a common issue with deep generative models, where they learn to generate an unrealistically reduced set of common outputs. We define diversity as the probability that a schedule is unique in the synthetic sample.
@@ 6 Results
We first summarise and compare all models at a high level. The density estimation of two selected models, the Discrete CNN and Continuous RNN, is then evaluated in more detail in Section 7. Results are supplemented with various scenarios in Sections 8 and 9 using our preferred model, the Continuous RNN.
@@
Table 7: Evaluation Summary
@@
Sample Quality
Invalidity 0.000 0.000 0.050 0.057 0.030 0.018 prob.
Creativity
Homogeneity 0.468 0.507 0.931 0.153 0.348 0.022 prob.
Conservatism 0.100 0.118 0.351 0.011 0.010 0.012 prob.
@@ 6.1 Density Estimation
Table 8: Density Estimation Evaluations
@@
Figures 12 and 12 present both aggregate activity frequencies (participations binned at 10-minute intervals) and disaggregate activity sequences, for the modelled and target real samples. Synthetic distributions should look like the target real distribution. Notably, we can see that the Discrete CNN has excellent aggregate performance, but is less convincing in disaggregate, for example, introducing additional stay-at-home sequences. The introduction of such sequences is common amongst all the discrete models.
@@
Figure 11: Aggregate Activity Frequencies (10-minute time bins)
Figure 12: Activity Sequences
@@ 6.2 Sample Quality
Table 9: Sample Quality Evaluations
@@ 6.3 Creativity
Table 10: Creativity Evaluations
@@ 7.1 Participations
The comparison of pair participations, which consider joint dependencies within schedules (anywhere within a schedule, not necessarily consecutive), is less clear. However, as per the aggregated results in Table 8 the Continuous RNN is better performing overall (an expected EMD of 0.004 versus 0.011). This is due to the Continuous RNN performing well across all activity combinations, including infrequent ones.
@@
Table 11: Activity Participation Rates - Mean Densities and EMD
@@ 7.2 Transitions
Table 12 details mean densities of transition rates for the ten most and ten least frequently observed bi-grams, tri-grams and quad-grams. These represent progressively longer sequences of transitions. The mean density of a transition can be interpreted as the expected number of each transition in a schedule. For example, the most common single transition, the bi-gram home-other has an average rate of 0.450 in the real sample. This can be interpreted as the expected number of such transitions in a schedule. The synthetic Discrete CNN and Continuous RNN samples have expected home-other rates of 0.414 and 0.403, underestimating the mean density of these transitions by only 0.04 of a single transition.
@@
Figure 13 shows the frequencies of whole schedule activity sequences. Both models maintain the pattern of common shorter sequences and uncommon longer sequences. However, the Discrete CNN introduces a significant increase in stay-at-home sequences, from 1.6% in the real sample to 8.2%. Conversely, the Continuous RNN under-generates stay-at-home sequences by about 50% (approximately 0.8% of schedules generated by the Continuous RNN are stay-at-home sequences).
Figure 13: Whole Schedule Sequences
Table 12: Activity Transition Rates - Mean Densities and EMD
@@ 7.3 Timing
Figure 14 compares the distributions of start times, end times and durations for different activities. Figure 15 compares the joint distributions of activity start times and durations. In both cases, the synthetic samples should resemble the target real sample. Notably, the Discrete CNN performs well at escort activity timing, capturing the bi-modal distribution well, but poorly at work activity timing, often generating unrealistically short work activities.
@@
As per Table 8, the Continuous RNN is better performing overall, however from Tables 13 and 14 we observe that the Discrete CNN tends to be better at the timing of short duration activities such as other, shop, visit, and medical, but worse at longer duration activities such as home, work, and education.
Figure 14: Activity Start Times and Durations
Figure 15: Joint Activity Start Times, End Times and Durations
Table 13: Activity Start Times and Durations
@@
Table 14: Joint Activity Start Times and Durations
@@
8 Sample Size Scenarios

### v3

@@ 4.2.1 Discrete Embedding and Un-Embedding
Un-embedding reverses the operation using soft-max to output activity probabilities. When generating schedules, we sample the highest probability activity at each time step, making the decoder architecture deterministic. Note that in all cases, we find sampling the highest probabilities at each step outperforms weighted sampling.
@@ 4.2.2 Continuous Embedding and Un-Embedding
The continuous encoding is un-embedded by the decoder following a similar process in reverse. A vector of durations for the sequence is separated and passed through a sigmoid activation so that all duration values are between zero and one. The remaining vectors are decoded into activity probabilities using a soft-max layer.
When generating schedules, we sample the highest probability activity at each time step, making the decoder architectures deterministic. The resulting activities and their durations for each sequence are then concatenated back together. After discarding padding tokens, durations for each sequence are normalised such that they sum to one.
@@ 4.3 Reconstruction Losses
Models using the discrete embedding use activity prediction categorical cross-entropy ($CE_{acts}$) as schedule reconstruction loss. Cross-entropy is calculated between activity prediction $p(a)$ and the ground truth $q(a)$, for each activity type in $\mathcal{A}$, at each step $n$. These are then averaged across all $N$ steps. The final discrete loss, $L_{discrete}$, including $D_{KL}$ is therefore:
$L_{discrete}=CE_{acts}+\beta D_{KL}.$ (6)
Where:
$CE_{acts}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{a\in\mathcal{A}}p(a_{n})\log q(a_{n})$ (7)
Models using the continuous encoding add an additional duration loss component to the reconstruction loss above. We use the mean squared error ($MSE_{durations}$) between the predicted duration $\hat{d}$ and ground truth duration $d$, for each schedule step $n$. This creates a continuous loss function, $L_{continuous}$. We introduce an additional hyperparameter $\alpha$, detailed in Table 6, to weight the duration loss component.
$L_{continuous}=CE_{acts}+\alpha MSE_{durations}+\beta D_{KL}.$ (8)
Where:
$MSE_{durations}=\frac{1}{N}\sum_{n=1}^{N}(d_{n}-\hat{d_{n}})^{2}$ (9)
@@ 4.4 Encoder/Decoder Block Architectures
The following section summarises the Feed-Forward, Convolutional and Recurrent VAE architectures. The number of blocks ($N$) and their size ($S$) vary for each model and are summarised in Table 6. The full architectures are available from Caveat [^4] https://github.com/big-ucl/caveat.
@@ 4.4.3 Recurrent (RNN) VAEs
The Recurrent (RNN) model architecture is presented in Figure  9. The encoder layer uses $N$ stacked Long-Short Term Memory (LSTM) units of size $N$ (Hochreiter and Schmidhuber, 1997). The block iteratively progresses through the input embeddings, and each LSTM unit passes a learnt hidden state as input to the next. The final output hidden state is then flattened as required for the latent block.
@@ 4.5 Model Training and Hyperparameters
Models are trained on 90% of the real sample data. We use the remaining 10% for validation during training. We train models until validation loss stabilises, typically for around 100 epochs. We do not find models suffer from over-fitting, likely due to the VAE design, which restricts information passing via the small latent layer and adds a stochastic process to training. We use Adam for gradient descent. Model training hyperparameters are reported in Table 6. The hyperparameters $\alpha$ and $\beta$ are chosen based on an extensive grid search to optimise VAE performances based on the evaluation criteria. The remaining hyperparameters are chosen to optimise validation loss using the Tree-structured Parzen Estimator by Optuna (Akiba et al., 2019). We therefore do not expect any significant model improvements to be possible via further hyperparameter exploration.
Table 6: Summary of VAE Model Hyperparameters
@@ 5 Evaluation Methodology
In this section, we consider the evaluation of synthetic samples of activity schedules generated by the various models. Each model is used to generate a sample of synthetic schedules $X_{syn}$, of equal size to the real sample $X_{real}$.
We aim for the generated or estimated distribution of schedules in  $X_{syn}$ to equal the distribution of schedules in $X_{real}$. This represents the ability of a model to correctly estimate the probability of any possible activity schedule - commonly called density estimation.
As per the problem statement (Section 2.3), we refer to the distribution of $X_{real}$ as $P_{X}$, and $X_{syn}$ as $\hat{P}_{X}$. We estimate the similarity of $P_{X}$ and $\hat{P}_{X}$ by extracting marginal distributions from the realised samples $X_{real}$ and $X_{syn}$, then comparing these using a distance metric.
We choose marginal distributions with qualitative meaning such that we can assess the implications of the quality of density estimation on specific tasks. We consider many marginal distributions such that when distances are aggregated, we can make a reasonable quantitative evaluation of overall density estimation. Density estimation is detailed in Section 5.1.
Perfect density estimation would satisfy our requirement for both individual schedule quality and aggregate representation across multiple schedules. However, because we can only estimate densities and because of the difficulty of directly comparing high-dimensional density estimates, we supplement evaluation by additionally considering metrics for individual schedule validity (Section 5.2) and creativity (Section 5.3). This is also consistent with generative evaluation from other domains reviewed in Section 1.5.
@@ 5.1.1 Marginal Distributions
Both participation and transition distributions are measured as rates, which are counts of occurrences for each schedule. We then consider the distribution of such rates across the whole sample.
@@
Figure 10 shows example marginal distributions extracted from the UK NTS. Rates are ordinal, and timings are continuous. However, the time distributions are binned at five-minute intervals to reduce computational complexity. All distributions can be described using their mean density, which can be interpreted as the expected rate or time. Expected values (mean densities) are shown in Figure 10 as vertical dashed lines.
@@ 5.1.2 Distances
Distances between the real and estimated distributions are measured using Earth Movers Distance (EMD). EMD is symmetric with meaningful values, such that we can reasonably interpret them. For example, a distance of 0.1 for work activity participation rate can be interpreted as an expectation of 0.1 fewer or more of work activities occurring in each schedule. Similarly, a distance of 0.5 for an activity start can be interpreted as that activity starting, on average, half a day late or early. For bivariate distributions, we use L1 distance (also known as Manhattan/city-block distance).
Figure 10: Example UK National Travel Survey Marginal Distributions
@@
Domain Distribution Segmentation Unit Descriptive Metric
@@ 5.3 Creativity Evaluation
We are additionally interested in the model’s ability to generate both (i) a good diversity of schedules – i.e. different from one another, as well as (ii) novelschedules – i.e. unique from schedules in the real sample. Diversity can be considered as a type of density evaluation metric, but is also used to specifically check if the model is suffering from mode collapse, a common issue with deep generative models, where they learn to generate an unrealistically reduced set of common outputs. We define diversity as the probability that a schedule is unique in the synthetic sample. Because diversity is not invariant to sample size, we keep all sample sizes consistent.
@@ 6 Results
We first summarise and compare all models at a high level. The density estimation of two selected models, the Discrete CNN and Continuous RNN, is then evaluated in more detail in Section 7. Results are supplemented with various scenarios in Sections 8.1 and 8.2 using our preferred model, the Continuous RNN.
@@
Table 9: Models Evaluation Summary: Domain-level distances, such that lower is better
@@
$\pm$ 0.008 $\pm$ 0.005 $\pm$ 0.054 $\pm$ 0.001 $\pm$ 0.001 $\pm$ 0.009
Validity 0.000 0.000 0.050 0.057 0.030 0.018 prob.
$\pm$ 0.001 $\pm$ $0$$\pm$ 0.079 $\pm$ 0.014 $\pm$ 0.006 $\pm$ 0.007
Creativity 0.284 0.313 0.641 0.082 0.179 0.017 prob.
$\pm$ 0.026 $\pm$ 0.011 $\pm$ 0.076 $\pm$ 0.012 $\pm$ 0.016 $\pm$ 0.003
(All results are means and standard deviations calculated from 5 model runs)
@@ 6.1 Density Estimation
Table 10: Models Density Estimation Evaluation: Distribution-level distances, such that lower is better
@@
(All results are means calculated from 5 model runs)
Figures 12 and 12 present both aggregate activity frequencies (participations binned at 10-minute intervals) and disaggregate activity sequences, for the modelled and target real samples. Estimated distributions should look like the target real distribution. Notably, we can see that the Discrete CNN has excellent aggregate performance, but is less convincing in disaggregate, for example, introducing additional stay-at-home sequences. The introduction of such sequences is common amongst all the discrete models.
@@
We speculate that the discrete models learn better aggregate patterns because time is consistently represented at each step within a sequence. This likely biases the models towards learning the most likely activity at each time step. In contrast, the continuous models cannot know the start time of each activity in a sequence without considering other predictions within the sequence. This likely biases the models towards better disaggregate sequences, but makes aggregate representation, particularly of timing, more challenging.
The DiscRNN performs the worst. This is likely because it is both (i) biased towards most likely activity estimation by the discrete encoding, and (ii) biased away from learning disaggregate sequence patterns due to having to pass a hidden state a relatively long way through the RNN architecture.
Figure 11: Comparison of Real and Synthesised Activity Frequencies: Aggregated activity participations in 10-minute time bins
Figure 12: Comparison of Real and Synthesised Activity Sequences: Disaggregate sequences of activity types
@@ 6.2 Sample Quality
Table 11: Model Validity Evaluations: Probabilities of being invalid, such that lower is better
@@
(All results are means calculated from 5 model runs)
@@ 6.3 Creativity
Table 12: Model Creativity Evaluations: Probabilities of being non-unique (homogeneity) or non-novel (conservatism), such that lower is better
@@
(All results are means calculated from 5 model runs)
6.4 Training and Validation Losses
Table 13 shows model training and validation losses. Corresponding training and validation losses are similar for each model, showing that they are not overfit. Model reconstruction losses (CE and MSE) are fairly consistent with final evaluations. Of the discrete models, the DiscRNN underperforms the DiscFF and DiscCNN. Of the continuous models, the ContRNN outperforms the ContFF and ContCNN, but only at duration prediction, suggesting the RNN architecture provides a better structure for duration reconstruction. Note that activity losses between the discrete and continuous models are not directly comparable.
Relatively high $D_{KL}$ is generally associated with relatively lower reconstruction losses, and vice versa. This is because where models struggle with reconstruction, they can instead minimise loss by minimising the $D_{KL}$ component. For example, the DiscRNN is particularly bad at reconstruction and hence has the lowest $D_{KL}$. The continuous models partly avoid this trade-off. For example, the ContCNN has both the lowest activity loss and $D_{KL}$ (of the continuous models). This suggests that the continuous representation is more suited to this problem. Perhaps because the continuous representation of time maps better to the continuous latent space. Further work might explore models with discrete latent spaces, such as the Vector Quantised-VAE by van den Oord et al. (2017).
Note that we use $\beta$ to regularise the impact of $D_{KL}$ on model learning as per Section 4.3. $\beta$ is carefully chosen as per Table 6 to optimise generative evaluation. During training, $\beta$ balances the tendency of the model to correctly distribute the latent layer or to reconstruct input schedules. Too low, and models excel at reconstruction, but generation evaluation is poor because the latent space is poorly distributed. Too high and models fail to reconstruct, leading to poor reconstruction losses and generation evaluation.
Table 13: Model Training and Validation Losses Summary
Discrete Continuous
FF CNN RNN FF CNN RNN
Training Set Losses
Activity Loss (CE) 0.106 0.1123 0.330 0.035 0.033 0.036
$\pm$ 0.046 $\pm$ 0.019 $\pm$ 0.081 $\pm$ 0.005 $\pm$ 0.005 $\pm$ 0.007
Duration Loss (MSE) - - - 0.002 0.002 0.000
- - - $\pm$ $0$$\pm$ $0$$\pm$ 0.000
$D_{KL}$ 11.274 8.041 3.364 8.366 5.792 8.194
$\pm$ 1.175 $\pm$ 0.394 $\pm$ 0.991 $\pm$ 0.222 $\pm$ 0.131 $\pm$ 0.123
Validation Set Losses
Activity Loss (CE) 0.106 0.112 0.329 0.034 0.033 0.036
$\pm$ 0.045 $\pm$ 0.018 $\pm$ 0.080 $\pm$ 0.006 $\pm$ 0.005 $\pm$ 0.006
Duration Loss (MSE) - - - 0.002 0.002 0.000
- - - $\pm$ $0$$\pm$ $0$$\pm$ 0.000
$D_{KL}$ 11.243 8.038 3.352 8.374 5.797 8.191
$\pm$ 1.185 $\pm$ 0.382 $\pm$ 0.984 $\pm$ 0.225 $\pm$ 0.133 $\pm$ 0.169
(All results are means and standard deviations from 5 model runs)
@@ 7.1 Participations
The comparison of pair participations, which consider joint dependencies within schedules (anywhere within a schedule, not necessarily consecutive), is less clear. However, as per the aggregated results in Table 10, the Continuous RNN is better performing overall (an expected EMD of 0.004 versus 0.011). An independent samples t-test additionally suggests a significant difference, t(5) = -8.101, p ¡ 0.001. This is due to the Continuous RNN performing well across all activity combinations, including infrequent ones.
@@
Table 14: DiscCNN and ContRNN Activity Participation Rates: Descriptions (mean densities) and Distances (EMD)
@@ 7.2 Transitions
Table 15 details mean densities of transition rates for the ten most and ten least frequently observed bi-grams, tri-grams and quad-grams. These represent progressively longer sequences of transitions. The mean density of a transition can be interpreted as the expected number of each transition in a schedule. For example, the most common single transition, the bi-gram home-other has an average rate of 0.450 in the real sample. This can be interpreted as the expected number of such transitions in a schedule. The synthetic DiscCNN and ContRNN samples have expected home-other rates of 0.414 and 0.403, underestimating the mean density of these transitions by only 0.04 of a single transition.
@@
Figure 12 shows the frequencies of whole schedule activity sequences. Both the DiscCNN and ContRNN maintain the pattern of common shorter sequences and uncommon longer sequences. However, the Discrete CNN introduces a significant increase in stay-at-home sequences, from 1.6% in the real sample to 8.2%. Conversely, the ContRNN under-generates stay-at-home sequences by about 50% (approximately 0.8% of schedules generated by the ContRNN are stay-at-home sequences).
Table 15: DiscCNN and ContRNN Activity Transition Rates: Descriptions (mean densities) and Distances (EMD)
@@ 7.3 Timing
Figure 14 compares the distributions of start times, end times and durations for different activities. Figure 14 compares the joint distributions of activity start times and durations. In both cases, the synthetic samples should resemble the target real sample.
Some activities, such as shop, education and escort, have more structured timing. These activities tend to take place at the same times of day, regardless of what else is in the activity sequence. The DiscCNN performs well at start and end times for these more structured activities due to its discrete representation, which ensures consistent activity timing at each step of a schedule. We consider the timing of these activities as more sequence invariant, because they are less affected by other activities in the schedule, but more by temporal constraints such as facility opening (and closing) times, which are consistent across all schedules.
However, the DiscCNN is biased towards shorter durations, often generating unrealistic short durations for longer activities such as work and education. This is because long-duration activities using the discrete representation are composed of many consecutive time steps. This increases the chance of an incorrect prediction. The result of an incorrect prediction is that an activity duration is shortened into two shorter duration activities. Where activities are typically shorter, this is less of an issue, so the DiscCNN performs very well at the timing of escort activities, which are both sequence invariant and of shorter duration.
In contrast to the DiscCNN, the ContRNN uses a continuous representation, which provides an explicit representation of activity durations, which is also directly included in the continuous loss function. This results in the ContRNN making better duration predictions than the Discrete CNN, regardless of activity length.
However, the continuous representation means that time is no longer persistent at each prediction step. Instead, it depends on both prior and later duration predictions. This makes the correct prediction of start and end times harder. Therefore, the Continuous RNN is less accurate at reconstructing activity start and end times. This is particularly notable for the more structured sequence invariant activities, such as escort.
@@
As per Table 10, the ContRNN is better performing overall, however from Tables 16 and 17 we observe that the Discrete CNN tends to be better at the timing of short duration activities such as other, shop, visit, and medical, but worse at longer duration activities such as home, work, and education.
Figure 13: DiscCNN and ContRNN Activity Start Times and Durations
Figure 14: DiscCNN and ContRNN Joint Activity Start Times, End Times and Durations
Table 16: DiscCNN and ContRNN Activity Start Times and Durations: Descriptions (mean densities) and Distances (EMD)
@@
** Mean distance metric
Table 17: DiscCNN and ContRNN Joint Activity Start Times and Durations: Descriptions (mean densities) and Distances (EMD)
@@
** Mean distance metric
7.4 Density Estimation Summary
As per Table 9, we find the ContRNN to outperform the DiscRNN at all domain-level distances. We formally confirm the significance of the improvement using independent sample Welch’s t-tests. We find strong evidence of improvement for participation density estimation ($p<0.001$) and weak evidence for transitions and timing ($p<0.1$).
8 Scenarios and Benchmarking
8.1 Sample Size Scenarios